# Can Frontier LLMs Replace Annotators in Biomedical Text Mining? Analyzing Challenges and Exploring Solutions

### v1

@@ Abstract
Large language models (LLMs) can perform various natural language processing (NLP) tasks through in-context learning without relying on supervised data. However, multiple previous studies have reported suboptimal performance of LLMs in biological text mining. By analyzing failure patterns in these evaluations, we identified three primary challenges for LLMs in biomedical corpora: (1) LLMs fail to learn implicit dataset-specific nuances from supervised data, (2) The common formatting requirements of discriminative tasks limit the reasoning capabilities of LLMs particularly for LLMs that lack test-time compute, and (3) LLMs struggle to adhere to annotation guidelines and match exact schemas, which hinders their ability to understand detailed annotation requirements which is essential in biomedical annotation workflow. To address these challenges, we experimented with prompt engineering techniques targeted to the above issues, and developed a pipeline that dynamically extracts instructions from annotation guidelines. Our findings show that frontier LLMs can approach or surpass the performance of state-of-the-art (SOTA) BERT-based models with minimal reliance on manually annotated data and without fine-tuning. Furthermore, we performed model distillation on a closed-source LLM, demonstrating that a BERT model trained exclusively on synthetic data annotated by LLMs can also achieve a practical performance. Based on these results, we explored the feasibility of partially replacing manual annotation with LLMs in production scenarios for biomedical text mining.
@@ 1 Introduction
1.1 Background: Annotation Paradigms in Biomedical Text Mining
Scientific literature contains a wealth of valuable information, presenting both opportunities and challenges for biological and medical research. Extracting knowledge from vast amounts of unstructured biomedical literature is critical for building biological databases, enabling researchers to uncover new insights and conduct data-driven studies. For example, the Gene Ontology (GO) project builds a comprehensive database by extracting information about genes and protein functions from scientific literature through a combination of manual curation and automated tools[1, 2]. This ontology database provides detailed classifications of biological processes, molecular functions, and cellular components of genes. Meanwhile, the ever-growing volume of published biomedical literature has exacerbated the issue of information overload, making it increasingly difficult for scientists, medical professionals, and the general public to quickly and accurately retrieve required information. For example, since the onset of the COVID-19 pandemic in 2020, over 400,000 papers related to COVID-19 and SARS-CoV-2 have been published in at least 8000 academic journals[3]. This urgents the need for effective automated text mining tools.
Extracting information from biological literature relies heavily on annotation conducted by human experts. Typically, curators define the annotation tasks and objectives, then collaborate with domain experts to develop annotation guidelines. Subsequently, human annotators conduct the annotation based on these guidelines[4, 5]. In many production use cases, manual annotation remains essential for continuously updating data to maintain annotation accuracy. Examples include projects like UniProtKB/Swiss-Prot and the Gene Ontology[6]. The labor-intensive nature of manual annotation has driven the adoption of natural language processing (NLP) methods in biomedical text mining[7]. These methods enable the automated extraction of valuable information from large volumes of unstructured text, including tasks such as identifying key entities (Named Entity Recognition, NER) and uncovering relationships between them (Relation Extraction, RE). Additionally, text classification plays a critical role in categorizing and organizing this information effectively. A well-known example of automated text mining in biological domain is PubTator, which applies machine learning techniques to annotate entities and relationships in biological literature, thereby assisting researchers in efficiently accessing and interpreting information[8, 9]. Another example is LitCovid, a web system for tracking the latest publications and research on COVID-19, which uses machine learning to classify newly published documents automatically[10, 11, 12]. However, even in such automated annotation workflows, human annotator are also required for building annotated datasets used for supervised learning models.
In the broader field of biomedical natural language processing (BioNLP), generative tasks such as medical question answering and text summarization also play a notable role. Nevertheless, the present study primarily focuses on information extraction, a category often referred to as discriminative tasks. Our focus on these tasks stems from the observation that large language models continue to face challenges in this domain, as will be discussed in detail in Section 1.2
@@ 1.1 Background: Annotation Paradigms in Biomedical Text Mining
1.2 Background: Studies on Language Models in BioNLP
Early biomedical text mining relied primarily on rule-based methods, statistical techniques, and early machine learning approaches[13, 14, 15, 16, 16]. In recent years, language models based on neural networks have achieved dominant performance: BERT based on Transformer encoder architecture have shown significant advances by leveraging pre-training on unannotated corpora to understand natural language and contextual knowledge[17, 18, 19]. These models can then be fine-tuned on smaller supervised datasets to perform various specific biomedical tasks. Despite their strong performance across diverse biomedical text mining applications, BERT-like models face a critical limitation that they depend on supervised data. As shown in Figure 1 b., the size of the corpus and the availability of labeled data directly influence the performance of these models[20].
The emergence of large language models (LLMs), exemplified by ChatGPT, seems to offer a potential solution to this challenge. LLMs are typically based on Transformer decoder architecture and are pre-trained on large-scale copora using next-token prediction to develop text generation capabilities[21][22][23]. Additionally, LLMs undergo alignment training like instruction-tuning to improve their ability to follow human instructions, engage in logical reasoning, and adapt to diverse tasks without requiring supervised data or specialized training[24].
As early as 2022, studies began applying models like GPT-3 to information extraction, showing promising results in clinical domain[25]. Following the release of ChatGPT, Qijie Chen et al. benchmarked GPT-3.5 on various tasks in the biomedical literature[26]. While the model outperformed baselines in Q&A tasks, its performance on information extraction tasks lagged significantly behind those of fine-tuned models. Similarly, a comprehensive evaluation of LLMs across 26 datasets conducted by Jahan et al. confirmed that these models struggle to match the performance of fine-tuned BERTs in discriminative tasks[27]. Further studies, such as those by Qingyu Chen et al. highlighted that even advanced proprietary models like GPT-4 which excel in biomedical generative tasks, exhibit notable performance deficits compared to fine-tuned BERT models in discriminative tasks[28]. In some datasets, GPT-4 achieved only 40% of the SOTA performance (e.g., 0.33 compared to 0.79). They furthermore analyzed the errors made by LLMs, but those analysis primarily focused on output format issues, without delving into the reasoning processes of the models.
@@ 1.3 Contribution of Our Work
We focus on three representative tasks in biomedical text mining—named entity recognition (NER), relation extraction (RE), and multi-label text classification. We first reproduced the previous evaluations of LLM in biological domain to analyze the failure patterns, then introduced series of prompt engineering methods and a information retrieval pipeline specially targeted those issues we identified:
(1) We revealed a significant discrepancy between the distribution of LLM outputs and the ground truth label across various dimensions such as the number and distribution of output labels, average span length of entity mentions. In contrast, models fine-tuned through supervised learning exhibit much smaller discrepancies. We assume this difference arises because supervised training implicitly captures dataset-specific features. To address this issue, we introduced an automatic prompt optimization techinique targeted to narrowing the gap between LLM-predicted labels and the correct labels.
@@
(3) Our analysis of LLM outputs highlights that many failures stem from a lack of understanding of task requirements. Annotation guidelines, which are often several pages to tens of pages long, specify crucial aspects such as task requirements, annotation categories, hierarchies, and granularities. They also provide examples, counterexamples, and instructions for handling edge cases[4, 5]. These details that are critical for accurate annotation but were largely overlooked in prior evaluations of LLMs. Thus, we explored incorporating annotation guidelines into prompts to address the issue of LLMs lacking an inherent understanding of task-specific requirements. Given that annotation guidelines are often lengthy, simply appending them to the prompt resulted in suboptimal performance, as LLM struggles to identify the most relevant sections. Baed on that, we manually divided the annotation guidelines into smaller text chunks and developed a Retrieval-Augmented Pipeline to dynamically retrieve the most relevant sections based on the task context.
@@
Utilizing the above approach, we achieved significant improvements in Biological Text Mining of LLM performance, surpassing the SOTA on 2 datasets and demonstrating practical performance levels on the remaining.
@@
Though some studies have shown that fine-tuning (including instruction tuning) may also be an effective approach[32, 33], the primary reason we did not adopt a fine-tuning approach is the substantial computational resources required for fine-tuning and deploying a large language model. Moreover, Nori et al. demonstrated that advanced prompt engineering can significantly enhance the performance of LLM and enabled a general-domain LLM to outperform domain-specific models[34]. Some reviews have also highlighted similar trends in related studies[35, 36, 37]. Additionally, inference-side scaling laws have garnered significant attention recently[38, 39, 40]. For example, GPT-o1 has shown exceptional performance in complex reasoning tasks by automatically generating reasoning steps before producing outputs[41].
@@
In real applications, deploying a LLM presents challenges. LLM demands significant GPU memory, and the attention mechanism used in GPT architectures has a quadratic order of memory and computation complexity with respect to the sequence length, which results in high inference costs[17, 21]. Cloud-based solutions, while feasible, incur considerable financial costs and pose potential data security risks. To mitigate these challenges, we adopted a workflow inspired by production use scenarios: human-annotated data is used to train machine learning models, which then continuously perform inference. We treated the LLM together with its output parser as a discriminative model and distilled it into a lighter model like BERT by a style of weak supervision learning. For the HoC and BC5CDR-RE datasets (which are a multi-label text classification dataset and a relation extraction dataset, respectively), we collected literature published after April 2024 following the same procedure as in the original studies[42, 43]. We then used the LLM to generate annotated data, which were used for fine-tuning a BERT model. The distilled BERT model achieved performance close to the previous state-of-the-art (SOTA).
@@
Our results suggest that for emerging text mining demands, relying solely on LLM inference in production environments or using an LLM+BERT distillation pipeline may be a pratical replacement of the widely adopted workflow of human annotation followed by machine learning.
@@ 2.1 Biomedical Tasks and Datasets
We focus on three key tasks: named entity recognition (NER), relation extraction (RE), and multi-label text classification (CLS). Biomedical named entity recognition aims to extract biomedical named entities such as genes, sequences, drugs, and treatments from text. For example, "Mutational analysis of CYP2C8 in hypertension using a new technology." In this case, we would identify: "CYP2C8" as "Gene", "hypertensive" as "Disease". Biomedical relation extraction involves identifying relationships between biomedical entities, such as drug-drug interactions. This task typically requires extracting (Entity 1, Entity 2, Relationship Type) triples. For example, in "The combination of warfarin and aspirin increases the risk of bleeding.", expected outputs include (Warfarin, Aspirin, Adverse Effect). Biomedical multi-label text classification involves assigning a document (sentence or article) to one or more labels. For instance, the possible labels of "Metformin is effective in controlling blood glucose levels and is also associated with gastrointestinal side effects." could be Effectiveness, Adverse Effect.
@@
We adopt evaluation metrics that were commonly adopted in prior research. For multi-label text classification and relation extraction, evaluation metrics include Macro Precision, Macro Recall, and Macro F1-score. For named entity recognition task, the evaluation is conducted using entity-level metrics. Their definitions are as follows:
  • Macro-Level Metrics: Macro Precision, Macro Recall, and Macro F1-score provide a class-agnostic evaluation of model performance by averaging the metrics across all classes without weighting by the number of instances per class. This ensures that all classes, regardless of size, contribute equally to the evaluation. Let $C$ denote the total number of classes:
    $\text{Precision}=\frac{1}{C}\sum_{i=1}^{C}\frac{\text{True Positives}_{i}}{\text{True Positives}_{i}+\text{False Positives}_{i}}$
    $\text{Recall}=\frac{1}{C}\sum_{i=1}^{C}\frac{\text{True Positives}_{i}}{\text{True Positives}_{i}+\text{False Negatives}_{i}}$
[…]
  • Entity-Level Metrics: Metrics are computed by aggregating predictions and ground truths for all entities across the entire dataset. Let CorrectEntities be the number of correctly predicted named entities, PredictedEntities be the total number of predicted entities, and GoldEntities be the total number of gold-standard (actual) entities. Then:
    $\text{Precision}=\frac{\text{CorrectEntities}}{\text{PredictedEntities}},\quad\text{Recall}=\frac{\text{CorrectEntities}}{\text{GoldEntities}},\quad\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@
BC5CDR-RE[43] 500/500/500 BioLinkBERT-Large, BioGPT Y (8 Pages)
@@ 2.2 Challenges Faced by LLM
In prior studies evaluating LLMs for biomedical text mining, the prompts typically included brief descriptions of the task type, labels, and output format. Building on these works, we reproduced their approaches using GPT-4 (API version: 0613[49]), consistent with Qingyu Chen’s research[28]) under a random 3-shot setting. The detailed results are presented in Table 2. Our evaluation scores are comparable to or slightly higher than, previous studies, which may be attributed to subtle differences in prompt text and improved robustness of the output parsers. For datasets that were previously only evaluated with GPT-3.5 or other LLMs, we observed improved results, likely due to the superior performance of GPT-4 itself[26, 27]. However, GPT-4 still significantly underperforms BERT-based models across all datasets except BC5CDR-RE.
To investigate the characteristics of outputs from LLMs, we analyzed the following statistics:
  • For named entity recolonization, we examined the number of entities per test instance and the average word length of each entity across the whole test set.
  • For relation extraction, we evaluated the number of relationships per test instance and the distribution of relationship categories.
  • For multi-label text classification, we analyzed the number of labels per test instance and the distribution of each label.
We selected four representative results to display in Figure 2, while additional detailed results are provided in Figures S1. From these results, we observe that across various feature metrics, BERT-based models generally align more closely with the ground truth labels in both the training and test sets compared to the outputs of GPT-4. Notably, GPT-4 achieved performance comparable to the fine-tuned state-of-the-art model (BioGPT) and outperformed fine-tuned BioBERT on the BC5CDR-RE dataset. However, the distribution of relations extracted by GPT-4 remained further from the test set distribution (one-sided Mann-Whitney U test, p-value = 9.25e-05, Wasserstein distance = 0.356) compared to BioBERT, which exhibited a closer match to the test set distribution (Wasserstein distance = 0.180).
Considering the reported superior performance of advanced closed-source models like GPT-4 in other generative tasks in biomedical domain and reports suggesting that their training data likely includes extensive biomedical literature[50, 51, 52, 53, 54], it is plausible that these models possess robust capabilities in understanding biomedical documents. This includes a strong grasp of domain-specific terminology, ambiguous expressions, and biomedical abbreviations. Given this, we hypothesize that the shortcomings of large language models (LLMs) in biomedical text mining tasks may partly stem from insufficient understanding of the datasets themselves. To validate this hypothesis, we conducted a case study focusing on the failures of LLMs in these tasks. These failures can be categorized as follows: 1. Errors that show some degree of reasonableness but fail to meet the detailed requirements about annotation categories, hierarchies, and granularities, or represent exceptions explicitly noted in the guidelines. 2. Errors arising from insufficient semantic understanding. 3. Parsing errors.
@@
1. Semantically Plausible but Inconsistent with Annotation Guidelines In many instances, predictions generated by GPT-4 exhibit a degree of rationality and may initially appear accurate. However, these predictions fail to conform to the specific and detailed requirements stipulated by the annotation guidelines, resulting in misalignment with the true labels. This category of errors can be further subdivided into three subtypes, though the boundaries between these subtypes are blurred: (a) Incorrect Annotation Hierarchy.
In the NCBI-Disease dataset, both mentions of specific diseases and disease classes (families of specific diseases) should be annotated. However, GPT-4 using prompt same as prior research often neglected disease class entities. For instance, in the sentence: “The RB1 gene mutation was investigated in a child with ectopic intracranial retinoblastoma using DNA obtained from both the pineal and retinal tumours of the patient.”(PMID: 9400934[55]) the LLM properly annotated all other disease mentions but failed to label “pineal and retinal tumours” as a disease class. In the BioRED (RE) dataset, only specific relationships (e.g., disease–chemical) among five defined entity types should be annotated. Nevertheless, the LLM labeled “dexmedetomidine” and “patient” as having an “Association” relationship in: “A comparison of severe hemodynamic disturbances between dexmedetomidine and propofol for sedation in neurocritical care patients.”(PMID: 24717468[56]) Where according to the guidelines, “patient” should be marked as a “Species” entity, which should not be linked in any relationship with a chemical entity in this dataset context.
@@
In BioRED (RE) dataset, relationships must be classified into categories like “positive correlation,” “negative correlation,” or “association.” However, the LLM sometimes misinterpreted the direction of effect. Consider this sentence: “…intravenous lidocaine infusion or a high dose of mexiletine …controlled the ventricular tachycardia.”(PMID: 15485686[57]) The LLM labeled the pair (mexiletine, ventricular tachycardia) as “positive correlation.” While it is true that mexiletine has a beneficial therapeutic effect on this disease, the guidelines specify that when a drug is used to treat a disease or reduces its severity, it should be labeled as “negative correlation”.
@@
According to the NCBI-Disease guidelines, words like “disease”, “syndrome”, “deficiency”, or “complication” alone should not be annotated. However, in the following text: “Low levels of beta hexosaminidase A in healthy individuals with apparent deficiency of this enzyme.”(PMID: 941901[58]) the LLM incorrectly annotated “deficiency” as a disease despite the guideline explicitly prohibiting such labeling. For the following sentence in BioRED (RE) dataset: “The acute toxicity of OPs is the result of their irreversible binding with AChEs in …”(PMID: 15036754[59]) LLM annotated “OP” and “toxicity” as having a “negative correlation,” even though the annotation guideline requires not to annotate the chemicals with the disease concept-“toxicity” directly, but to annotate the disease due to the “toxicity” and “drug-induced” instead.
2. Errors due to Inadequate Semantic Comprehension These are pure reasoning mistakes by the LLM and are not necessarily tied to annotation guidelines. For instance, in the following article snippet from BC5CDR (RE) dataset: “Late-onset scleroderma renal crisis induced by tacrolimus and prednisolone: a case report. Scleroderma renal crisis (SRC) is a rare complication of systemic sclerosis (SSc) but can be severe enough to require temporary or permanent renal replacement therapy. Moderate to high dose corticosteroid use is recognized as a major risk factor for SRC. Furthermore, there have been reports of thrombotic microangiopathy precipitated by cyclosporine in patients with SSc.”(PMID:22836123[60]) The correct chemical–disease interactions (CDI) that should be extracted include: Cyclosporine–microangiopathy, Corticosteroid–SRC/SSc, tacrolimus–SRC/SSc. However, the LLM omitted (tacrolimus, SRC/SSc) and misidentified the disease associated with “cyclosporine” as SRC/SSc instead of microangiopathy. Interestingly, in a simpler prompt setting (e.g., “retrieve all the CID pairs from the text”), the LLM produced the correct pairs in a more free-text style. This suggests that forcing the LLM to output fully structured annotations may disrupt its internal reasoning processes. This observation motivated the technique we propose in Section 3.1.3.
3. Parsing Errors A parsing error occurs when a large language model (LLM) fails to produce output in the specified structured format, rendering the output parser unable to process the result correctly. While models like LLaMA2-13B have frequently reported these inconsistencies according to Q. Chen et al[28, 61], we have not encountered such errors outside the development phase. This can be attributed to that models like GPT-4 benefit from extensive training focused on generating structured outputs, as well as that we designed a robust output parser capable of effectively capturing edge cases.

### v2

@@ Abstract
Multiple previous studies have reported suboptimal performance of LLMs in biomedical text mining. By analyzing failure patterns in these evaluations, we identified three primary challenges for LLMs in biomedical corpora: (1) LLMs fail to learn implicit dataset-specific nuances from supervised data, (2) The common formatting requirements of discriminative tasks limit the reasoning capabilities of LLMs particularly for LLMs that lack test-time compute, and (3) LLMs struggle to adhere to annotation guidelines and match exact schemas, which hinders their ability to understand detailed annotation requirements which is essential in biomedical annotation workflow. We experimented with prompt engineering techniques targeted to the above issues, and developed a pipeline that dynamically extracts instructions from annotation guidelines. Our results show that frontier LLMs can approach or surpass the performance of state-of-the-art (SOTA) BERT-based models with minimal reliance on manually annotated data and without fine-tuning. Furthermore, we performed model distillation on a closed-source LLM, demonstrating that a BERT model trained exclusively on synthetic data annotated by LLMs can also achieve a practical performance. Based on these findings, we explored the feasibility of partially replacing manual annotation with LLMs in production scenarios for biomedical text mining.
@@ 1 Introduction
1.1 Background: Biomedical Text Mining and the Role of Language Models
Biomedical literature contains valuable knowledge essential for building curated biological databases and enabling data-driven discovery. Information extraction from this literature has traditionally relied on manual annotation, where domain experts develop annotation guidelines, and human annotators label data accordingly. This labor-intensive process remains indispensable in high-precision scenarios, such as maintaining resources like UniProtKB/Swiss-Prot and the Gene Ontology databases[2].
The labor-intensive nature of manual annotation has driven the adoption of automated biomedical text mining approaches—such as named entity recognition (NER), relation extraction (RE), and multi-label classification[7]. Tools like PubTator[9] and LitCovid[11] demonstrate the feasibility of applying machine learning to assist with information extraction and document triaging. Complementing this trend is the rise of BERT-based models, which have become the mainstream approach in recent years[18]. BERT usually undergone self-supervised pretraining and are then fine-tuned on task-specific supervised datasets to perform diverse biomedical text mining tasks. However, these methods remain fundamentally dependent on manually annotated data (Figure 1).
@@ 1.1 Background: Biomedical Text Mining and the Role of Language Models
The emergence of large language models (LLMs) like ChatGPT offers a potential paradigm shift. LLMs are typically based on Transformer decoder architecture and are pre-trained on large-scale copora using next-token prediction to develop text generation capabilities[23]. Additionally, LLMs undergo alignment training to improve their ability to follow human instructions, engage in logical reasoning, and adapt to diverse tasks without specialized training.
As early as 2022, studies began applying models like GPT-3 to information extraction, showing promising results in clinical domain[25]. Following the release of ChatGPT, Qijie Chen et al. benchmarked GPT-3.5 on various tasks in the biomedical literature[26]. While the model outperformed baselines in Q&A tasks, its performance on information extraction tasks lagged significantly behind those of fine-tuned models. Similarly, a comprehensive evaluation of LLMs across 26 datasets conducted by Jahan et al. confirmed that these models struggle to match the performance of fine-tuned BERTs in discriminative tasks[27]. Further studies, such as those by Qingyu Chen et al. highlighted that even advanced proprietary models like GPT-4 which excel in biomedical generative tasks, exhibit notable performance deficits compared to fine-tuned BERT models in discriminative tasks[28]. In some datasets, GPT-4 achieved only 40% of the SOTA performance. They furthermore analyzed the errors made by LLMs, but those analysis primarily focused on output format issues, without delving into the reasoning processes of the models. Overall, those studies revised that despite advancements in LLMs, fine-tuned BERT models remain superior for biomedical text mining tasks, highlighting a gap that still needs to be addressed.
@@ 1.2 Contribution of Our Work
We target three core biomedical text mining tasks: named entity recognition (NER), relation extraction (RE), and multi-label text classification. NER extracts biomedical entities (e.g., genes, drugs, diseases) from text. RE identifies relationships among these entities, such as drug–disease interactions. Multi-label classification assigns multiple relevant labels to biomedical texts at the sentence or document level.
We first reproduced the previous evaluations of LLM in biomedical domain to analyze specific failure patterns, then introduced series of prompt engineering methods and a information retrieval pipeline specially targeted those issues we identified:
(1) We revealed a significant discrepancy between the distribution of LLM outputs and the ground truth label across various dimensions such as the number and distribution of output labels, average span length of entity mentions. We assume this difference arises because supervised training implicitly captures dataset-specific features. To address this, we introduced an automatic prompt optimization technique targeted to find a dataset-specific prompt narrowing the gap between LLM-predicted labels and the correct labels.
@@
(3) Many LLM failures were traced to insufficient understanding of task-specific annotation guidelines, which are often several pages to tens of pages long, specify crucial aspects such as task requirements, annotation categories, hierarchies, granularities, and instructions for handling edge cases. These details that are critical for accurate annotation but were largely overlooked in prior evaluations of LLMs. However, since annotation guidelines are often lengthy, LLMs struggle to identify the most relevant sections when these guidelines are directly appended to the prompt, resulting in suboptimal performance. To address this, we manually divided the annotation guidelines into smaller text chunks and developed a retrieval-augmented pipeline to dynamically retrieve the most relevant sections based on task context.
@@
Utilizing the above approach, we achieved significant improvements in biomedical text mining of LLM performance, surpassing the SOTA on 2 datasets and demonstrating practical performance levels on the remaining.
@@
Though some studies have shown that fine-tuning (including instruction tuning) may also be an effective approach[32, 33], the primary reason we did not adopt a fine-tuning approach is the substantial computational resources required for fine-tuning and deploying a large language model. Moreover, it is widely recognized that advanced prompt engineering can substantially improve the performance of LLMs, enabling general-domain models to achieve or even surpass domain-specific models[34].
@@
Additionally, deploying LLMs is highly challenging, and relying on cloud services introduces significant financial costs and potential data security risks. To address this, we adopted a workflow inspired by annotation pipelines in current production scenarios: human-annotated data is used to train machine learning models, which then continuously perform inference. We treated the LLM with its output parser as a discriminative model and distilled it into a lighter BERT-based model by weak supervision learning on two datasets. We collected literature following the same procedure as in the original studies and used the LLM to generate annotated data, which were used to fine-tune a BERT model. The distilled BERT model achieved performance close to the previous SOTA.
@@
Our results suggest that for emerging text mining demands, relying solely on LLM inference in production environments or using an LLM+BERT distillation pipeline may be a practicall replacement of the widely adopted workflow of human annotation followed by machine learning.
@@ 2.1 Biomedical Tasks and Datasets
BC5CDR-RE[43] 1323/1343/1375 BioLinkBERT-Large, BioGPT Y (8 Pages)
@@ 2.2 Challenges Faced by LLM
In prior studies evaluating LLMs for biomedical text mining, the prompts typically included brief descriptions of the task type, labels, and output format[28, 26, 27]. We reproduced their approaches using GPT-4 (API version: 0613[49]), following Qingyu Chen’s research) under a random 3-shot setting, and analyzed failure patterns of the outputs (Table 2). Our evaluation scores are comparable to or slightly higher than previous studies, which may be attributed to subtle differences in prompt text and improved robustness of the output parsers. However, LLM still significantly underperforms BERT-based models across all datasets except BC5CDR-RE.
We analyzed the following statistics for each tasks, to investigate the characteristics of outputs from LLMs:
  • Named entity recolonization: the number of entities per test instance and the average word length of each entity across the whole test set.
  • Relation extraction: the number of relationships per test instance and the distribution of relationship categories.
  • Multi-label text classification: the number of labels per test instance and the distribution of each label.
We selected four representative results shown in Figure 2, with additional details provided in Figures S1 and S2. Across various metrics, BERT-based models generally produced outputs closer to the ground truth labels compared to GPT-4. Notably, GPT-4 achieved performance comparable to the SOTA model (fine-tuned BioGPT) and outperformed fine-tuned BioBERT on the BC5CDR-RE dataset. However, the distribution of relations extracted by GPT-4 remained further from the test set distribution (one-sided Mann-Whitney U test, p-value = 9.25e-05, Wasserstein distance = 0.356) compared to BioBERT, which exhibited a closer match to the test set distribution (Wasserstein distance = 0.180).
Considering the reported superior performance of advanced closed-source models like GPT-4 in other generative tasks in biomedical domain and reports suggesting that their training data likely includes extensive biomedical literature[50, 54], we hypothesize that its shortcomings in biomedical text mining partially stem from insufficient understanding of dataset-specific requirements, rather than from limitations in understanding biomedical documents themselves. To validate this, we conducted a case study focusing on the failures of LLMs in these tasks (Table 2).
@@
1. Semantically Plausible but Inconsistent with Annotation Guidelines Predictions appear reasonable but fail to comply with the annotation guideline details, leading to incorrect labels. Examples include following: (a) Incorrect Annotation Hierarchy.
In the NCBI-Disease dataset, both mentions of specific diseases and disease classes (families of specific diseases) should be annotated. However, GPT-4 using prompt same as prior research sometimes neglected disease class entities. For example: “The RB1 gene mutation was investigated in a child with ectopic intracranial retinoblastoma using DNA obtained from both the pineal and retinal tumours of the patient.”(PMID: 9400934[55]) the LLM properly annotated all other disease mentions but failed to label “pineal and retinal tumours” as a disease class.
@@
In BioRED (RE) dataset, relationships must be classified into categories like “positive correlation,” “negative correlation,” or “association.” However, the LLM sometimes misinterpreted the direction of effect. For example: “…intravenous lidocaine infusion or a high dose of mexiletine …controlled the ventricular tachycardia.”(PMID: 15485686[57]) The LLM labeled the pair (mexiletine, ventricular tachycardia) as “positive correlation.” While it is true that mexiletine has a beneficial therapeutic effect on this disease, the guidelines specify that when a drug is used to treat a disease or reduces its severity, it should be labeled as “negative correlation”.
@@
According to the guideline of NCBI-Disease, words like “disease”, “syndrome”, or “deficiency” alone should not be annotated. However, in the following: “Low levels of beta hexosaminidase A in healthy individuals with apparent deficiency of this enzyme.”(PMID: 941901[58]) the LLM incorrectly annotated “deficiency” as a disease despite the guideline explicitly prohibiting that.
2. Errors due to Inadequate Semantic Comprehension Some errors are purely due to reasoning mistakes by the LLM and are independent of annotation guidelines. Examples are shown in Appendix Table S1. Interestingly, for the case in Table S1.2, the LLM produced correct answers in a free-text style when given a simpler prompt, instead of one requiring detailed structured outputs. This suggests that forcing structured outputs may disrupt the LLM’s internal reasoning. This observation motivated our proposed technique in Section 3.1.3.
3. Parsing Errors Parsing errors occur when an LLM’s output does not follow the required structured format, causing parsing failures. While models like LLaMA2-13B were reported to frequently exhibit such issues [28, 61], we did not encounter them outside the development stage. This can be attributed to that frontier LLMs benefit from extensive training focused on generating structured outputs, as well as that we implemented a robust output parser capable of capturing edge cases.